OpenStudy (anonymous):

Find the sum. 17x3 + (3x + 8x3) =

OpenStudy (anikhalder):

17x3 + (3x + 8x3) =51+3x+24 =75+3x =3(x+25)

OpenStudy (anonymous):

agreed ^

OpenStudy (anonymous):

17x^3 + (3x + 8x^3) = 17x^3+8x^3 +3x = 25x^3 + 3x

OpenStudy (anonymous):

@Jmichael did you right x cubed or x *3 ?

OpenStudy (anonymous):

*write

OpenStudy (anikhalder):

haha..you are probably right @tomo

OpenStudy (anonymous):

x cubed